I am being asked to "Find the linear functions satisfying": F(4)=3 F(-4)=11
I know the two problems should be set up as follows...
3=m(4)+b and 11=m(-4)+b
I also know that you have to find B first but, I don't know how to do this...

Answer by stanbon(75887)

"Find the linear functions satisfying": F(4)=3 F(-4)=11
---
You have two points: (4,3) and (-4,11)
---
slope = (11-3)/(-4-4) = 8/-8 = -1
---------------------
Form: y = mx+b
3 = -1(4) + b
b = 7
-------
Equation:
y = -x+7
==================
Cheers,
Stan H.
==================